# Neural Sign Actors: A diffusion model for 3D sign language production from text

### 05 Dec 2023 v1

@@ Abstract
Sign Languages (SL) serve as the predominant mode of communication for the Deaf and Hard of Hearing communities. The advent of deep learning has aided numerous methods in SL recognition and translation, achieving remarkable results. However, Sign Language Production (SLP) poses a challenge for the computer vision community as the motions generated must be realistic and have precise semantic meanings. Most SLP methods rely on 2D data, thus impeding their ability to attain a necessary level of realism. In this work, we propose a diffusion-based SLP model trained on a curated large-scale dataset of 4D signing avatars and their corresponding text transcripts. The proposed method can generate dynamic sequences of 3D avatars from an unconstrained domain of discourse using a diffusion process formed on a novel and anatomically informed graph neural network defined on the SMPL-X body skeleton. Through a series of quantitative and qualitative experiments, we show that the proposed method considerably outperforms previous methods of SLP. We believe that this work presents an important and necessary step towards realistic neural sign avatars, bridging the communication gap between Deaf and hearing communities. The code, method and generated data will be made publicly available.
@@ 1 Introduction
Sign language is a form of language in which visual-manual modalities are used instead of spoken words to convey meaning. It is the predominant form of communication for more than 70 million Deaf and Hard of Hearing people around the world. Akin to verbal languages, sign languages have extremely rich vocabulary and grammar, yet the complexities differ drastically [53]. To enable effective visual communication, sign languages consist of both manual and non-manual components [35]. The manual modality encompasses hand articulation, orientation, position, and motion, while non-manual elements include arm movements and facial expressions [6]. Whilst it is possible to convey some meaning using just hand articulations, expressiveness is limited since non-manual elements often convey emotions [53, 3].
With the advent of deep learning, several methods have been proposed to bridge the domain gap between sign and spoken languages. Most methods focus on Sign Language Recognition (SLR) which includes the translation of a specific sign to its corresponding meaning, as well as Sign Language Translation (SLT) that extends SLR to the translation of a sign sequence to its spoken word equivalent. This is usually tackled using glosses [10, 28, 14, 15], which are simplified mid-level representations that relate each sign with a corresponding meaning. However, even though glosses have provided a substantial enhancement to SLT methods, they have a predefined informative bottleneck, which limits the translation accuracy, and they usually fail to provide long-range dependencies and contextual information [10, 32, 62].
@@ 2.2 Sign Language Datasets
A major contributing factor to the slow-paced advancements in sign language research is the absence of large-scale datasets [8]. Earlier datasets were designed with a focus on sign language recognition using isolated signs [58, 4, 57, 31, 33], containing a limited vocabulary. To address the challenges of sign language recognition and translation within the context of complete sentences, several continuous sign language datasets have been introduced. More specifically, RWTH-BOSTON-50 [59], Dreuw et al. [18], SIGNUM [1], and BSL [50] along with DictaSign Corpus, which was developed in several languages [7, 20, 21], were among the first datasets with sentence level annotations. While additional datasets featuring an expanded set of signs have been introduced [12, 31, 33], it is crucial to emphasize the importance of continuous sign language for the purposes of translation and production. S-pot [56] was among the first large-scale continuous sign language datasets with over 1K signs of Finnish sign language collected in a constrained environment. To enforce the robustness of sign language translation methods, RWTH-Phoenix [9] contained a collection of TV clips with German sign language, remaining amongst one of the most popular datasets used for sign language translation [10] and production [47]. Similarly, BSL-1K, as presented in [2], curated a dataset featuring British sign language (BSL) used in casual conversations, encompassing a total of 1K signs. Recently, Duarte et al. [19] collected How2Sign, a large-scale dataset of American Sign Language (ASL), that is aligned with speech signals from the How2 dataset [44]. How2Sign is equipped with over 16K signs, captured in a total of seventy-nine hours of continuous sign language. A major limitation of the above datasets, is the lack of available 3D annotations, that not only aid the translation tasks, but are also essential for training realistic 3D signing avatars. In this work, we have extended the How2Sign dataset by incorporating high-quality SMPL-X [41] annotations, thereby establishing it as the first publicly available 3D sign language dataset.
@@ 3 Dataset
where $\Pi_{K}$ is the intrinsic camera projection matrix. Following extensive experimentation, we observed that optimization is only necessary for the arm and hand joints, as the OSX-regressed body joints are sufficiently accurate. Fitting hand poses using 2D keypoints is an exceptionally challenging task, primarily due to the numerous articulations and the inherent ambiguities within the solutions. While several methods have been proposed to constrain SMPL to feasible body poses [41, 17, 55], pose prior models for the hand remain unexplored. To constrain the optimization to plausible human and hand poses, we propose a simple but intuitive approach using Principal Component Analysis (PCA) to model the subspace of anatomically feasible poses. Specifically, we trained a PCA pose prior model on two large datasets of human body [38] and hand [22] poses, to model the distribution of feasible arm and hand poses. To formulate the prior loss we measure the reconstruction error of a mesh $\mathbf{X}$ projected and reconstructed from the PCA space $\mathbf{U}$ as:
@@ 5 Experiments
Table 1: Mean per vertex error of the proposed and the baseline 3D reconstruction pipelines on the SGNify mocap dataset [25]. Reconstruction error is measured in mm.
@@ 5.1 Dataset Evaluation
Figure 4: Qualitative comparison between the proposed and the baseline fitting frameworks on SGNify [25] and How2Sign [19] datasets.
In Tab. 1, we report the reconstruction error of the proposed pipeline and compare it with OSX [36] and the SGNify method, which is the current state-of-the-art method for 3D fitting from SL videos. It must be noted that unlike the SGNify method, the proposed fitting pipeline achieves a smaller reconstruction error despite not including any SL-driven losses. Indeed, the SGNify pipeline fails to produce accurate reconstructions, while at the same time, it occasionally generates non-feasible articulations which can be qualitatively validated in Fig. 4. The powerful prior model that guides the fitting optimization leads our method to valid poses and articulation.
@@ 5.4 Ablation
The proposed method consists of four main components: the anatomically inspired pose encoder, the pose and expression embedding layers, the autoregressive decoder, and the text encoding. In this section, we evaluate the contribution of each component to the final generations of the model. For qualitative assessments, we refer the reader to the supplementary material.
Effect of Pose Encoder and Embedding Layers. Initially, we ablate the novel pose encoder and we substitute it with an MLP layer (w/o GNN), similar to the expression encoder. Under this setting, the generated motions are smoother and mainly deviate around the mean pose.
Aligned with our hypothesis, updating poses in an anatomically inspired manner enables the network to learn higher frequencies, such as rare articulations, without an increase in the network’s capacity.
Additionally, as mentioned in Sec. 4.2, in the traditional setting of motion diffusion models, during the denoising step, poses and expressions are sampled from a Gaussian distribution and are then decoded back to their original space using a permutation equivariant network, such as an MLP. Such permutation equivariance, treats poses under a uniform setting that limits the generative power of the network. As shown in Tab. 3, without pose and expression embedding layer (w/o Pose and Expression Embedding), the model fails to produce any reasonable sign, resulting in a performance drop under all metrics.
@@
Effect of Text Encoding. The text encoding module has a pivotal effect on motion generation. Firstly, we substitute the CLIP encoder with a word-level embedding layer that is trained with the rest of the method in an end-to-end fashion. We set the embedded size to 256 although we did not observe significant differences in performance. Following this, we utilized the pretrained DistilBERT [45], whose parameters remain frozen throughout training. As depicted in Tab. 3, training word embeddings from scratch strongly affects the generalization of the network, leading to large MPVPE and MPJPE metrics. In contrast, DistilBERT achieves better performance than learnable word embeddings, although it does not outperform CLIP embeddings. This is aligned with our assumption that sentence embeddings could provide better insights regarding the meaning of a sentence compared to word-level embeddings. Especially in the task of sign language production, where there is not an explicit one-to-one mapping between words and poses, sentence level embeddings provide a more powerful text encoding solution.
@@ 6 Conclusion
Neural 3D sign language production is an important challenge that aims to aid the Deaf and Hard of Hearing community and can effectively increase their inclusiveness in any social environment. In this work, we made a step towards high fidelity 3D SLP, by deriving a large-scale 3D dataset to train a text conditioned diffusion-based model. We initially introduce a precise 3D sign language reconstruction pipeline that outperforms previous SL reconstruction methods. Then, we train a motion generative model using an autoregressive diffusion model. The core of our method is founded on a novel, anatomically inspired, graph neural network that learns the pose distribution and enables highly detailed articulations. Importantly, leveraging the powerful CLIP text embeddings, the proposed model can generalize to out-of-distribution samples. Extensive experiments on sign language generation tasks, including a perceptual study with ASL fluent subjects, demonstrate the superiority of the proposed method compared to the previous approaches.

### 05 Apr 2024 v2

@@ Abstract
[…]
@@ 1 Introduction
Sign language (SL) is a form of language in which visual-manual modalities are used instead of spoken words to convey meaning. It is the predominant form of communication for more than 70 million Deaf and Hard of Hearing people around the world. Akin to verbal languages, SLs have extremely rich vocabulary and grammar, yet the complexities differ drastically [55]. To enable effective visual communication, they consist of both manual and non-manual components [35]. The manual modality encompasses hand articulation, orientation, position, and motion, while non-manual elements include arm movements and facial expressions [6]. Whilst it is possible to convey some meaning using just hand articulations, expressiveness is limited since non-manual elements often convey emotions [55, 3].
Recently, several methods have been proposed to bridge the domain gap between sign and spoken languages. Most methods focus on Sign Language Recognition (SLR) which includes the translation of a specific sign to its corresponding meaning, as well as Sign Language Translation (SLT) that extends SLR to the translation of a sign sequence to its spoken word equivalent. This is usually tackled using glosses [10, 28, 14, 15], which are simplified mid-level representations that relate each sign with a corresponding meaning. However, even though glosses have provided a substantial enhancement to SLT methods, they have a predefined informative bottleneck, which limits the translation accuracy, and they usually fail to provide long-range dependencies and contextual information [10, 32, 64].
@@ 2.2 Sign Language Datasets
A major contributing factor to the slow-paced advancements in sign language research is the absence of large-scale datasets [8]. Earlier datasets were designed with a focus on sign language recognition using isolated signs [60, 4, 59, 31, 33], containing a limited vocabulary. To address the challenges of sign language recognition and translation within the context of complete sentences, several continuous sign language datasets have been introduced. More specifically, RWTH-BOSTON-50 [61], Dreuw et al. [18], SIGNUM [1], and BSL [52] along with DictaSign Corpus, which was developed in several languages [7, 20, 21], were among the first datasets with sentence level annotations. While additional datasets featuring an expanded set of signs have been introduced [12, 31, 33], it is crucial to emphasize the importance of continuous sign language for the purposes of translation and production. S-pot [58] was among the first large-scale continuous sign language datasets with a vocabulary of over 1K signs of Finnish sign language collected in a constrained environment. To enforce the robustness of sign language translation methods, RWTH-Phoenix [9] contained a collection of TV clips with German sign language, remaining amongst one of the most popular datasets used for sign language translation [10] and production [49]. Similarly, BSL-1K, as presented in [2], curated a dataset featuring British sign language (BSL) used in casual conversations, encompassing a total vocabulary of 1K signs. Recently, Duarte et al. [19] collected How2Sign, a large-scale dataset of American Sign Language (ASL), that is aligned with speech signals from the How2 dataset [46]. How2Sign is equipped with a vocabulary of over 16K signs, captured in a total of seventy-nine hours of continuous sign language. A major limitation of the above datasets, is the lack of available 3D annotations, that not only aid the translation tasks, but are also essential for training realistic 3D signing avatars. In this work, we have extended the How2Sign dataset by incorporating high-quality SMPL-X [41] annotations, thereby establishing it as the first publicly available 3D sign language dataset.
@@ 3 Dataset
where $\Pi_{K}$ is the intrinsic camera projection matrix. Following extensive experimentation, we observed that optimization is only necessary for the arm and hand joints, as the OSX-regressed body joints are sufficiently accurate. Fitting hand poses using 2D keypoints is an exceptionally challenging task, primarily due to the numerous articulations and the inherent ambiguities within the solutions. While several methods have been proposed to constrain SMPL to feasible body poses [41, 17, 57], pose prior models for the hand models [44, 42] remain unexplored. To constrain the optimization to plausible human and hand poses, we propose a simple but intuitive approach using Principal Component Analysis (PCA) to model the subspace of anatomically feasible poses. Specifically, we trained a PCA pose prior model on two large datasets of human body [38] and hand [22] poses, to model the distribution of feasible arm and hand poses. To formulate the prior loss we measure the reconstruction error of a mesh $\mathbf{X}$ projected and reconstructed from the PCA space $\mathbf{U}$ as:
@@ 5 Experiments
Table 1: Mean per vertex error (mm) of the proposed and the baseline methods on the SGNify mocap dataset [25].
@@ 5.1 Dataset Evaluation
Figure 4: Qualitative comparison between the proposed and the baseline fitting frameworks on SGNify [25] and How2Sign [19].
In Tab. 1, we report the reconstruction error of the proposed pipeline and compare it with OSX [36] and the SGNify method, which is the current state-of-the-art method for 3D fitting from SL videos. It must be noted that unlike the SGNify method, the proposed fitting pipeline achieves a smaller reconstruction error despite not including any SL-driven losses. The powerful prior model that guides the fitting optimization leads our method to valid poses and articulation Fig. 4.
@@ 5.4 Ablation
The proposed method consists of four main components: the anatomically inspired pose encoder, the pose and expression embeddings, the autoregressive decoder, and the text encoding. In this section, we evaluate the contribution of each component to the final generations of the model.
Effect of Pose Encoder and Embedding Layers. Initially, we ablate the novel pose encoder and we substitute it with an MLP layer (w/o GNN), similar to the expression encoder. Under this setting, the generated motions are smoother and mainly deviate around the mean pose. Aligned with our hypothesis, updating poses in an anatomically inspired manner enables the network to learn higher frequencies, such as rare articulations, without an increase in the network’s capacity. Additionally, as mentioned in Sec. 4.2, in the traditional setting of motion diffusion models, during the denoising step, poses and expressions are sampled from a Gaussian distribution and are then decoded back to their original space using a permutation equivariant network, such as an MLP. Such permutation equivariance, treats poses under a uniform setting that limits the generative power of the network. As shown in Tab. 3, without pose and expression embedding layer (w/o Pose and Expression Embedding), the model fails to produce any reasonable sign, resulting in a performance drop under all metrics.
@@
Effect of Text Encoding. The text encoding module has a pivotal effect on motion generation. Firstly, we substitute the CLIP encoder with a word-level embedding layer that is trained with the rest of the method in an end-to-end fashion. We set the embedded size to 256 although we did not observe significant differences in performance. Following this, we utilized the pretrained DistilBERT [47], whose parameters remain frozen throughout training. As depicted in Tab. 3, training word embeddings from scratch strongly affects the generalization of the network, leading to large MPVPE and MPJPE metrics. In contrast, DistilBERT achieves better performance than learnable word embeddings, although it does not outperform CLIP embeddings. This is aligned with our assumption that sentence embeddings could provide better insights regarding the meaning of a sentence compared to word-level embeddings. Especially in the task of SLP, where there is not an explicit one-to-one mapping between words and poses, sentence level embeddings provide a more powerful text encoding solution.
@@ 6 Conclusion
Neural 3D sign language production is an important challenge that aims to aid the Deaf and Hard of Hearing community and can effectively increase their inclusion in any social environment. In this work, we made a step towards high fidelity 3D SLP, by deriving a large-scale 3D dataset to train a text conditioned diffusion-based model. The release of additional relevant databases will enable the training of even more robust architectures. We initially introduce a precise 3D sign language reconstruction pipeline that outperforms previous SL reconstruction methods. Then, we train a motion generative model using an autoregressive diffusion model. The core of our method is founded on a novel, anatomically inspired, graph neural network that learns the pose distribution and enables highly detailed articulations. Importantly, leveraging the powerful CLIP text embeddings, the proposed model can generalize to out-of-distribution samples. Extensive experiments on sign language generation tasks, including a perceptual study with ASL fluent subjects, demonstrate the superiority of the proposed method compared to the previous approaches.
Acknowledgements. S. Zafeiriou was supported by EPSRC Project DEFORM (EP/S010203/1) and GNOMON (EP/X011364). R.A. Potamias was supported by EPSRC Project GNOMON (EP/X011364).